OpenStudy (anonymous):

assume the probability that you will make a sale on any given telephone call is 0.19. Find the probability that u (a) make your first sale on the fifth call(b)make your scale on the first second, or thisd call and(c) do not make a scal on teh first three calls I know the answer to part (a)... please only help me with pasrt (b) and (c)

OpenStudy (anonymous):

It is a binomial distribution what you have to use here: \[P(k)=\left(\begin{matrix}n \\ k\end{matrix}\right)p^k(1-p)^{n-k}\] Case b) n=3, k=1,2,3 \[P=\left(\begin{matrix}3 \\ 1\end{matrix}\right)0.19^1\times 0.81^2+\left(\begin{matrix}3 \\ 2\end{matrix}\right)0.19^2\times 0.81^1+\left(\begin{matrix}3 \\ 3\end{matrix}\right)0.19^3\times 0.81^0\] Case c) n=3, k=0 \[P=\left(\begin{matrix}3 \\ 0\end{matrix}\right)0.19^0\times 0.81^3\] Does it make sense to you?